# An Insight into Security Code Review with LLMs: Capabilities, Obstacles and Influential Factors

### v3

@@ 3.5.1. Attributes Construction
First, we collected the attributes from previous works (Paul et al., 2021; McIntosh et al., 2016; Thongtanunam et al., 2017) that were utilized in analyzing relationships related to code review and applicable to the code file data in our work. Given that LLMs understand and generate text token by token, we adjusted the factor adopted from McIntoshet al.’s work (McIntosh et al., 2016) — ‘Size’, to ‘Token’ (see the first row of Table 3), as larger code files typically contain more tokens. Next, we extracted characteristics of data that were explicitly indicated to affect or correlate with the capabilities of LLMs in Chen et al.’s work  (Chen et al., 2023), transforming them into factors for our experiment. Lastly, after thoroughly observing and analyzing the distribution of responses, we further created and adopted several potentially influential factors based on our observations. In the end, we obtained a list of 10 factors in Table 3, along with the source, definition, and rationale for each factor.
@@
Since acquiring a factor like ‘AuthorExperience’ entails querying the code review history of projects, we utilized Gerrit API to retrieve detailed information for all patchsets under each code change. We used Python scripts complemented with manual analysis to query, compute, and label these factors. Besides, since we adopted ‘SecurityDefectType’ in our factor list, we only considered data where the code file contains a single security defect to avoid the case in which ‘SecurityDefectType’ could have multiple values. After filtering out code files with multiple security defects, we aggregated the dataset from three repetitive experiments. We obtained a total of 1,215 responses generated by the GPT-4 and $\textbf{P}_{cid}$ combination. This comprised 405 responses from each experiment and their factor values ready for model fitting.
3.5.2. Regression Model Construction
@@
Correlation and Redundancy Analysis: If explanatory variables are highly correlated, it can cause interference in model construction and analysis. As the predefined explanatory variables in our study include both categorical and continuous variables, we utilized a refined Pearson’s hypothesis test of independence — Phi_K (Baak et al., 2020), to measure the correlation between each factor. The correlation coefficients obtained are illustrated in Fig. 3. We chose $|\rho|>0.7$ as the threshold because it is recommended as the threshold for strong correlation (Hinkle et al., 2003). From Fig. 3, where an abbreviation denotes each factor, it can be seen that the correlation coefficients between each pair of factors are all below 0.7, suggesting that there is no strong correlation between these factors. However, two factors without high correlation can still be duplicates, distorting the relationship between explanatory and response variables in model fitting. We encoded categorical variables as dummy variables for redundancy analysis (Legendre and Legendre, 2012). Utilizing the redun function in the rms package, we assessed potential redundant variables with its default threshold of $R^{2}\geq 0.9$, ultimately finding none.
Figure 3. The Phi_K correlation coefficients of explanatory variables
@@
Degrees of Freedom Allocation: To assign degrees of freedom to each explanatory variable, we measured the potential of nonlinear relationships between each factor and the response variable by calculating the Spearman’s rank correlations (${\rho}^{2}$) between them. As Harrell  (Harrell, 2001) recommended, explanatory variables with larger ${\rho}^{2}$ can be assigned more degrees of freedom within the range of 3 to 5. According to the computation results shown in Fig. 4, for interval variables, we considered using the rcs function from the rms package to allocate 3 degrees of freedom for the factor ‘Token’ with the highest ${\rho}^{2}$, while the other factors are allocated 1 degree of freedom. Category variables should be converted into factor variables, and their corresponding degrees of freedom are automatically allocated during model fitting.
@@
Figure 4. Dotplot of the Spearman ${\rho}^{2}$ between explanatory and response variables across three repetitive experiments
@@
All the results and scripts utilized in data labelling, extraction, and regression analysis are provided in our replication package (Yu et al., 2024).
4. Results
4.1. Performance of LLMs (RQ1)
Based on the evaluation metrics defined in Section 3.3.3, we evaluated the capability of six popular LLMs in security code review across five different prompts. We compared the performance of LLMs with baseline tools in the C/C++ and Python datasets, respectively. If two of the three metrics were superior (i.e., higher I-Score or IH-Score, lower M-Score), we deemed that the current combination outperforms the others. Due to differences in the token limit of each LLM and the inherent token count of each prompt template, the number of responses successfully generated by each LLM-prompt combination also varied. To ensure rigorous comparisons, we selected the intersection of source code files in our complete dataset so that all LLM-prompt combinations could generate responses, resulting in 57 instances. As a result, two sets of scores are provided: (1) the response count and performance scores of each LLM-prompt combination on the complete Python and C/C++ dataset (see Table 4 and Table 5), and (2) the performance scores of each LLM-prompt combination on the selected 57 instances (see Table 6). The best values for each score are highlighted in bold and marked with asterisks in the tables.
From Table 4 and Table 5, it is evident that, on the Python dataset, the combination of GPT-4 with $\textbf{P}_{cid}$ achieved the best performance. However, on the C/C++ dataset, the combination of Llama 2 70B with $\textbf{P}_{cid}$ performs the best. However, Table 6 demonstrated that in the selected 57 cases, Llama 2 70B scored lower than GPT-4. We speculate that this discrepancy could be attributed to the smaller volume of responses obtained by Llama 2 70B, leading to inflated performance scores of Llama 2 70B in the complete datasets. On the whole, we consider GPT-4 paired with $\textbf{P}_{cid}$ as the top-performing LLM-prompt combination. As of the time of our experiment (May 2024), the capabilities of currently popular LLMs in conducting security code reviews are still limited. Notably, with enhanced prompts, the best performance of each LLM in security code review is significantly superior to that of static analysis tools. This could be attributed to two reasons: (1) for context-sensitive tools such as CodeQL, static analysis can be impacted by the lack of code context in our dataset (Russo and Sabelfeld, 2010); (2) there is a relatively high frequency of security defects related to multi-threading and asynchronous programming, such as race conditions. As such features can only be accurately modelled using dynamic analysis, these security defects are challenging for static analysis tools to detect (Charoenwet et al., [n. d.]).
Table 4. Performance scores and entropy for each LLM-Prompt combination on the Python dataset
@@ 4.1. Performance of LLMs (RQ1)
Table 5. Performance scores and entropy for each LLM-Prompt combination on the C/C++ dataset
@@
Table 6. Performance scores for each LLM-Prompt combination on 57 instances that all LLM-prompt combinations could generate responses
@@
We can see from Table 4 and Table 5 that compared to using the basic prompt $\textbf{P}_{b}$, the performance of GPT-3.5 significantly decreases under $\textbf{P}_{c}$ and $\textbf{P}_{cid}$ in both Python and C/C++ datasets. The performances of GPT-4 with $\textbf{P}_{c}$ exhibit different changes across different datasets, with an improvement in the C/C++ dataset but a decrease in the Python dataset. However, all other LLMs achieved higher performance scores across both datasets. Moreover, GPT-4, GPT-4 Turbo, Gemini Pro, Llama 2 7B, and Llama 2 70B all achieved their best performance under $\textbf{P}_{cid}$. Therefore, it can be inferred that adding information related to CWE to the prompt generally enhances the capabilities of security defect detection of various LLMs in code review. It is worth noting that under $\textbf{P}_{cid}$, the decrease in GPT-4’s I-Score (Python: 4.76% →3.24%, C/C++: 4.45% →3.91%) and in M-Score (Python: 61.31% →46.47%, C/C++:55.47% →44.53%) reflect that guiding LLMs to reference CWE in the prompt indeed enables GPT-4 to detect more security defects. Still, it does not guarantee the accuracy of the detection results. Under $\textbf{P}_{cid}$, providing a specific CWE list as a reference in the prompt leads to significant improvements across all metrics and datasets for GPT-4. Hence, how to leverage CWE information in the prompt is worthy of consideration.
4.1.3. RQ1.3
@@
4.1.4. RQ1.4
According to the mean entropies of various LLM-prompt combinations in Table 4 and Table 5, GPT-3.5, GPT-4 Turbo, and Gemini Pro exhibit relatively higher consistency in their responses across three repetitive experiments, followed by Llama 2 7B and Llama 2 70B, with GPT-4 showing the least consistency. Specifically, on the Python dataset, the highest consistency of responses was observed under $\textbf{P}_{cot}$ with GPT-4 Turbo, with a mean entropy of 0.0404. On the C/C++ dataset, the combination of GPT-3.5 and $\textbf{P}_{c}$ demonstrates the lowest mean entropy of 0.0195. Given these two LLM-prompt combinations’ poor security defect detection capabilities, their high consistency may be due to their tendency not to detect any security defects in the code. GPT-4 paired with $\textbf{P}_{cid}$ is considered the top-performing combination in security code review but shows high mean entropies on both Python and C/C++ datasets, with 0.9416 and 0.6955 respectively. These results suggest that LLMs with more robust security defect detection capabilities may also exhibit greater randomness.
4.2. Quality Problems in Responses (RQ2)
The average distribution of 12 problem types across four themes in responses to three repetitive experiments is illustrated in Fig. 5, where the same color represents types within the same theme, with darker shades indicating the themes.
Figure 5. Average proportion of each problem type present in responses generated by GPT-4
From the perspective of the four themes, most responses contain problems related to Conciseness, accounting for 90.18%. Nearly half of the responses (41.56%) involve Compliance related problems, which indicate the requirements of the task described in the prompt are not met. Moreover, 27.08% of the responses show problems of Understandability due to inappropriate logic, expression, or attribution. Notably, problems related to Correctness, such as deviation from objective facts or common knowledge in the field of security, are encountered less frequently than other themes (14.08%). The typical types of problems within each theme are described in detail below.
@@ 4.2. Quality Problems in Responses (RQ2)
Understandability: Problems that reduce understandability of responses include Vague statement (21.38%), Incoherent expression (1.19%), Logical contradiction (1.11%), and Groundless speculation (5.26%). Among these, Vague statement is the most prevalent one. It refers to a situation where the description of the detected security defect is not specific enough, failing to pinpoint the defect to a specific location or providing a overly general fix suggestion. An example (Example 2) of Vague statement is shown below, in which GPT-4 did not provide a concrete location or line number for the identified security defect in the response.
@@
Conciseness: In all responses, 11.4% contain Redundant content, 15.12% include information Irrelevant to the task of security defect detection, and 86.38% contain Unnecessary information, all of which are problems related to conciseness of responses. Among them, Unnecessary information, which refers to information that is somewhat related to security code review but does not contribute to detecting security defects, frequently appears in the responses generated by GPT-4. Such information includes the reasons for the absence of security defects in the current code and general suggestions on security measures and review processes. As illustrated in the example below (Example 3), the response indicates that no defect related to CWE-284 exists in the code and further elaborates on the reason. However, this is not the content we required GPT-4 to generate in the prompt, hence is deemed to be Unnecessary information.
@@
Compliance: Three types of non-compliance were identified regarding whether the responses satisfied the requirements in the prompts. Among them, Detection failure is observed in very few responses (0.15%), where GPT-4 explicitly claimed that it could not perform security check in its response or misinterpreted requirements in the prompts for other tasks. Non-compliant output was present in 20.43% of responses, in which GPT-4 fails to include the fixed phrase ‘No security defects are detected in the code’, when no defects are found. Incomplete answer is the most common problem (20.98%) under this theme, referring to cases in which GPT-4’s responses lack the line numbers and fix suggestions of detected security defects requested in the prompt. From the example below (Example 4), we can see that while the potential memory management issues caused by temporary files and directories are described and a fix suggestion is provided, no specific line numbers are given. Hence, it is categorized as an Incomplete answer.
@@
While Kabir et al. (Kabir et al., 2024) also evaluated the quality problems in LLM-generated responses, the application domain of LLMs and the distribution of identified problems in their work differ from ours. Specifically, Kabir et al. requested LLMs to answer programming questions on Stack Overflow, finding that over half (52%) of the responses suffered from correctness issues. In contrast, our study applied LLMs to security code review, revealing a much lower incidence of correctness issues at 14.81%. This indicates that LLMs face different challenges across various software engineering tasks, highlighting the need to optimize and customize LLMs for specific areas of application.
4.3. Factors Influencing LLMs (RQ3)
4.3.1. Model Stability

### v4

@@ 2.5.1 Attributes Construction
First, we collected the attributes from previous works (Paul et al., 2021; McIntosh et al., 2016; Thongtanunam et al., 2017) that were utilized in analyzing relationships related to code review and applicable to the code file data in our work. Given that LLMs understand and generate text token by token, we adjusted the factor adopted from McIntosh et al. (2016) — ‘Size’, to ‘Token’ (see the first row of Table 3), as larger code files typically contain more tokens. Next, we extracted characteristics of data that were explicitly indicated to affect or correlate with the capabilities of LLMs in Chen et al. (2023), transforming them into factors for our experiment. Lastly, after thoroughly observing and analyzing the distribution of responses, we further created and adopted several potentially influential factors based on our observations. In the end, we obtained a list of 10 factors in Table 3, along with the source, definition, and rationale for each factor.
@@
Since acquiring a factor like ‘AuthorExperience’ entails querying the code review history of projects, we utilized Gerrit API to retrieve detailed information for all patchsets under each code change. We used Python scripts complemented with manual analysis to query, compute, and label these factors. Besides, since we adopted ‘SecurityDefectType’ in our factor list, we only considered data where the code file contains a single security defect to avoid the case in which ‘SecurityDefectType’ could have multiple values. We aggregated the dataset from three repetitive experiments after filtering out code files with multiple security defects. We obtained a total of 1,215 responses generated by the GPT-4 and $\textbf{P}_{cid}$ combination. This comprised 405 responses from each experiment, and their factor values were ready for model fitting.
2.5.2 Regression Model Construction
@@
Correlation and Redundancy Analysis: If explanatory variables are highly correlated, it can cause interference in model construction and analysis. As the predefined explanatory variables in our study include both categorical and continuous variables, we utilized a refined Pearson’s hypothesis test of independence — Phi_K (Baak et al., 2020), to measure the correlation between each factor. The correlation coefficients obtained are illustrated in Fig. 3. We chose $|\rho|>0.7$ as the threshold because it is recommended as the threshold for strong correlation (Hinkle et al., 2003). From Fig. 3, where an abbreviation denotes each factor, it can be seen that the correlation coefficients between each pair of factors are all below 0.7, suggesting that there is no strong correlation between these factors. However, two factors without high correlation can still be duplicates, distorting the relationship between explanatory and response variables in model fitting. We encoded categorical variables as dummy variables for redundancy analysis (Legendre and Legendre, 2012). Utilizing the redun function in the rms package [^2] https://cran.r-project.org/web/packages/rms/index.html, we assessed potential redundant variables with its default threshold of $R^{2}\geq 0.9$, ultimately finding none.
Figure 3: The Phi_K correlation coefficients of explanatory variables
@@
Degrees of Freedom Allocation: To assign degrees of freedom to each explanatory variable, we measured the potential of nonlinear relationships between each factor and the response variable by calculating Spearman’s rank correlations (${\rho}^{2}$) between them. As  Harrell (2001) recommended, explanatory variables with larger ${\rho}^{2}$ can be assigned higher degrees of freedom.  Paul et al. (2021) suggested limiting the maximum allocated degrees of freedom for an explanatory variable to five or fewer, thus reducing the risk of overfitting. According to the computation results shown in Fig. 4, for interval variables, we considered using the rcs function from the R’s rms package to allocate 3 degrees of freedom for the factor ‘Token’ with the highest ${\rho}^{2}$, while the other factors are allocated 1 degree of freedom. Category variables should be converted into factor variables, and their corresponding degrees of freedom are automatically allocated during model fitting.
@@
Figure 4: Dotplot of the Spearman ${\rho}^{2}$ between explanatory and response variables across three repetitive experiments
@@
All the results and scripts utilized in data labelling, extraction, and regression analysis are provided in our replication package (Yu et al., 2025).
3 Results
3.1 Performance of LLMs (RQ1)
Based on the evaluation metrics defined in Section 2.3.3, we evaluated the capability of six popular LLMs in security code review across five different prompts. We compared the performance of LLMs with baseline tools in the C/C++ and Python datasets, respectively. If two of the three metrics were superior (i.e., higher I-Score or IH-Score, lower M-Score), we deemed that the current combination outperformed the others. Due to differences in the token limit of each LLM and the inherent token count of each prompt template, the number of responses successfully generated by each LLM-prompt combination also varied. To ensure rigorous comparisons, we selected the intersection of source code files in our complete dataset so that all LLM-prompt combinations could generate responses, resulting in 57 instances. As a result, two sets of scores are provided: (1) the response count and performance scores of each LLM-prompt combination on the complete Python and C/C++ dataset (see Table 4 and Table 5), and (2) the performance scores of each LLM-prompt combination on the selected 57 instances (see Table 6). The best values for each score are highlighted in bold and marked with asterisks in the tables.
From Table 4 and Table 5, it is evident that, on the Python dataset, the combination of GPT-4 with $\textbf{P}_{cid}$ achieved the best performance. However, on the C/C++ dataset, the combination of Llama 2 70B with $\textbf{P}_{cid}$ performs the best. However, Table 6 demonstrated that in the selected 57 cases, Llama 2 70B scored lower than GPT-4. We speculate that this discrepancy could be attributed to the smaller volume of responses obtained by Llama 2 70B, leading to inflated performance scores of Llama 2 70B in the complete datasets. On the whole, we consider GPT-4 paired with $\textbf{P}_{cid}$ as the top-performing LLM-prompt combination. As of the time of our experiment (May 2024), the capabilities of currently popular LLMs in conducting security code reviews are still limited. Notably, with enhanced prompts, the best performance of each LLM in security code review is significantly superior to that of static analysis tools. This could be attributed to two reasons: (1) for context-sensitive tools such as CodeQL, static analysis can be impacted by the lack of code context in our dataset (Russo and Sabelfeld, 2010); (2) there is a relatively high frequency of security defects related to multi-threading and asynchronous programming, such as race conditions. As such, features can only be accurately modelled using dynamic analysis, and these security defects are challenging for static analysis tools to detect (Charoenwet et al., 2024).
Table 4: Performance scores and entropy for each LLM-Prompt combination on the Python dataset
@@ 3.1 Performance of LLMs (RQ1)
Table 5: Performance scores and entropy for each LLM-Prompt combination on the C/C++ dataset
@@
Table 6: Performance scores for each LLM-Prompt combination on 57 instances that all LLM-prompt combinations could generate responses
@@
We can see from Table 4 and Table 5 that compared to using the basic prompt $\textbf{P}_{b}$, the performance of GPT-3.5 significantly decreases under $\textbf{P}_{c}$ and $\textbf{P}_{cid}$ in both Python and C/C++ datasets. The performances of GPT-4 with $\textbf{P}_{c}$ exhibit different changes across different datasets, with an improvement in the C/C++ dataset but a decrease in the Python dataset. However, all other LLMs achieved higher performance scores across both datasets. Moreover, GPT-4, GPT-4 Turbo, Gemini Pro, Llama 2 7B, and Llama 2 70B all achieved their best performance under $\textbf{P}_{cid}$. Therefore, it can be inferred that adding information related to CWE to the prompt generally enhances the capabilities of security defect detection of various LLMs in code review. It is worth noting that under $\textbf{P}_{cid}$, the decrease in GPT-4’s I-Score (Python: 4.76% →3.24%, C/C++: 4.45% →3.91%) and in M-Score (Python: 61.31% →46.47%, C/C++:55.47% →44.53%) reflects that guiding LLMs to reference CWE in the prompt indeed enables GPT-4 to detect more security defects. Still, it does not guarantee the accuracy of the detection results. Under $\textbf{P}_{cid}$, providing a specific CWE list as a reference in the prompt leads to significant improvements across all metrics and datasets for GPT-4. Hence, leveraging CWE information in the prompt is worthy of consideration.
3.1.3 RQ1.3
@@
3.1.4 RQ1.4
According to the mean entropies of various LLM-prompt combinations in Table 4 and Table 5, GPT-3.5, GPT-4 Turbo, and Gemini Pro exhibit relatively higher consistency in their responses across three repetitive experiments, followed by Llama 2 7B and Llama 2 70B, with GPT-4 showing the least consistency. Specifically, on the Python dataset, the highest consistency of responses was observed under $\textbf{P}_{cot}$ with GPT-4 Turbo, with a mean entropy of 0.0404. On the C/C++ dataset, the combination of GPT-3.5 and $\textbf{P}_{c}$ demonstrates the lowest mean entropy of 0.0195. Given these two LLM-prompt combinations’ poor security defect detection capabilities, their high consistency may be due to their tendency not to detect any security defects in the code. GPT-4 paired with $\textbf{P}_{cid}$ is considered the top-performing combination in security code review but shows high mean entropies on both Python and C/C++ datasets, with 0.9416 and 0.6955 respectively. These results suggest that LLMs with more robust security defect detection capabilities may also exhibit greater non-determinism.
3.2 Quality Problems in Responses (RQ2)
3.2.1 RQ2.1
Figure 5: Average proportion of each problem type present in responses generated by GPT-4
The average distribution of 12 problem types across four themes in responses to three repetitive experiments is illustrated in Fig. 5, where the same color represents types within the same theme, with darker shades indicating the themes. From the perspective of the four themes, most responses contain problems related to Conciseness, accounting for 90.18%. Nearly half of the responses (41.56%) involve Compliance related problems, which indicate the requirements of the task described in the prompt are not met. Moreover, 27.08% of the responses show problems of Understandability due to inappropriate logic, expression, or attribution. Notably, problems related to Correctness, such as deviation from objective facts or common knowledge in the field of security, are encountered less frequently than other themes (14.08%). The typical types of problems within each theme are described in detail below.
@@ 3.2 Quality Problems in Responses (RQ2)
Understandability: Problems that reduce understandability of responses include Vague statement (21.38%), Incoherent expression (1.19%), Logical contradiction (1.11%), and Groundless speculation (5.26%). Among these, Vague statement is the most prevalent one. It refers to a situation where the description of the detected security defect is not specific enough, failing to pinpoint the defect to a specific location or providing an overly general fix suggestion. An example (Example 2) of Vague statement is shown below, in which GPT-4 did not provide a concrete location or line number for the identified security defect in the response.
@@
Conciseness: In all responses, 11.4% contain Redundant content, 15.12% include information Irrelevant to the task of security defect detection, and 86.38% contain Unnecessary information, all of which are problems related to conciseness of responses. Among them, Unnecessary information, which refers to information that is somewhat related to security code review but does not contribute to detecting security defects, frequently appears in the responses generated by GPT-4. Such information includes the reasons for the absence of security defects in the current code and general suggestions on security measures and review processes. As illustrated in the example below (Example 3), the response indicates that no defect related to CWE-284 exists in the code and further elaborates on the reason. However, this is not the content we required GPT-4 to generate in the prompt, hence it is deemed to be Unnecessary information.
@@
Compliance: Three types of non-compliance were identified regarding whether the responses satisfied the requirements in the prompts. Among them, Detection failure is observed in very few responses (0.15%), where GPT-4 explicitly claimed that it could not perform a security check in its response or misinterpreted requirements in the prompts for other tasks. Non-compliant output was present in 20.43% of responses, in which GPT-4 fails to include the fixed phrase ‘No security defects are detected in the code’, when no defects are found. Incomplete answer is the most common problem (20.98%) under this theme, referring to cases in which GPT-4’s responses lack the line numbers and fix suggestions of detected security defects requested in the prompt. The example below (Example 4) shows that while the potential memory management issues caused by temporary files and directories are described and a fix suggestion is provided, no specific line numbers are given. Hence, it is categorized as an Incomplete answer.
@@
While Kabir et al. (2024) also evaluated the quality problems in LLM-generated responses, the application domain of LLMs and the distribution of identified problems in their work differ from ours. Specifically, Kabir et al. (2024) requested LLMs to answer programming questions on Stack Overflow, finding that over half (52%) of the responses suffered from correctness issues. In contrast, our study applied LLMs to security code review, revealing a much lower incidence of correctness issues at 14.81%. This indicates that LLMs face different challenges across various software engineering tasks, highlighting the need to optimize and customize LLMs for specific application areas.
3.2.2 RQ2.2
We selected all items related to hallucination from the quality problem types in RQ1.1 and mapped them to the predefined set of hallucination subtypes. The details of mapping, corresponding explanations, and proportion results are shown in Table 7. Since a single response may exhibit multiple types of hallucination simultaneously, the sum of the proportions for each subtype exceeds the proportion of responses containing at least one hallucination. Overall, 20.19 % of the responses contain hallucinations. Among these, Factual Contradiction is the most frequent subtype, accounting for 10.13%, followed by Context Inconsistency and Factual Fabrication, which appear in 6.25% and 5.62% of responses, respectively. To alleviate these hallucinations, Retrieval-Augmented Generation (RAG) can be employed to incorporate external factual information (Tonmoy et al., 2024; Lewis et al., 2020), and post-training can also be applied to enhance the model’s capability to capture and retain contextual knowledge (Kumar et al., 2025).
Table 7: Mapping of hallucination and quality problem types with corresponding proportions.
Halluc. Type* Subtype QP Type* Explanation Prop.*
Factuality Hallucination Factual Contradiction Incorrect Concept In Incorrect Concept cases, the model’s output contradicts the authoritative definition of the corresponding CWE, thus categorized as Factual Contradiction. 10.13%
Factual Fabrication Groundless Speculation Groundless speculation refers to subjective assumptions made without contextual evidence or knowledge support. As it involves fabricated factual content, it is categorized as Factual Fabrication. 5.62%
Faithfulness Hallucination Instruction Inconsistency Detection Failure Detection Failure refers to cases where the model fails to perform the security checks or misinterprets the task, leading to deviations from the instruction. Therefore, it is categorized as Instruction Inconsistency. 0.15%
Context Inconsistency Incorrect Fact Incorrect Fact denotes cases where the line numbers, identifiers, or code content generated by the model do not align with the provided code context, thus categorized as Context Inconsistency. 6.25%
Logical Inconsistency Logical Contradiction Logical Contradiction refers to cases where the model’s reasoning about a security defect is internally contradictory, thus classified as Logical Inconsistency. 1.11%
Any Hallucination 20.19%
    Abbreviations: Halluc. = Hallucination, QP = Quality Problem, Prop. = Proportion.
3.3 Factors Influencing LLMs (RQ3)
3.3.1 Model Stability